Discussion Week 4

As the office manager of a large group practice you have two computational tasks to undertake today.

  1. You need to run a Depreciation Computation: Units-of-Service

Using the format in Table 9–1, compute Units-of-Service Depreciation using the following assumptions:

Cost to be depreciated = $50,000

Salvage value = zero

Total units of service = 10,000

Units of service per year: Year 1 = 2,200; Year 2 = 2,100; Year 3 = 2,300; Year 4 = 2,200; Year 5 = 200

  1. Next you need to calculate the FTEs needed next year to Annualize Staffing

To annualize your staffing plan, you need to convert the staff’s net paid days worked to a factor. The office is open and staffed seven days a week, due to your contracts with two managed care plans. The group practice pays for 8 holidays, 5 sick days, and 1 education day each year. An employee must work one full year to earn 5 vacation days. An employee must have worked full time for three full years before earning 10 annual vacation days. Because the turnover is so high, nobody on staff has earned more than 5 vacation days.

Required

   1. Compute Net Paid Days Worked for a full-time employee in the physicians’ group.

   2. Convert Net Paid Days Worked to a factor so you can annualize your staffing plan.
Answer #1
1
Year Book Value year start Units of activity Depreciation Exp. Acc. Dep Book Value year-end
1               50,000          2,200                12,222         12,222              37,778
2               37,778          2,100                11,667         23,889              26,111
3               26,111          2,300                12,778         36,667              13,333
4               13,333          2,200                12,222         48,889                1,111
5                 1,111             200                 1,111         50,000                      0
         9,000
Depreciation per Units = (50,000 - 0)/9,000 = 5.5555 per unit of activity
2
CATEGORY OF DAYS
Total days: business year 364
Less 2 days off per week 104
Number of paid days per year 260
Less: Paid days not worked
Holidays 8
Sick days 5
Education days 1
Vacation days 5
Total non-productive days 19
1 Net productive days 241
2 Factor = 364/241 = 1.51037
(Total days in the business year divided by net paid days worked)